# Failure of the measure contraction property via quotients in higher-step sub-Riemannian structures

### v1

@@ 2.1. Measure contraction property
Remark 2.3 (Variants).
Definition 2.1 should be compared with the one-sided measure contraction property of Sturm, introduced independently, see [Stu06b, Rmk. 5.11 $(iii)$]. A variant in terms of the Rényi entropy was used by Cavalletti and Mondino in [CM17] (see also [CM21, Def. 6.8]). From [Raj12, Sec. 5], one can deduce that the variant in [CM17] implies Definition 2.1, see [CM21, Lemma 6.13]. Since we do not assume the essentially non-branching condition, these variants may not be equivalent. Nevertheless, with minor modifications in the proofs, any variant would have led to the same conclusion for what concerns the results of this paper.
Any variant of $\mathrm{MCP}$ implies a corresponding Bishop-Gromov volume growth inequality for $(X,\mathsf{d},\mathsf{m})$, and all its consequences. In particular, we record that a metric measure space satisfying the $\mathrm{MCP}(K,N)$ with $K>0$ (and $N>1$) must be compact, and that setting
$A_{t,x}:=\left\{\gamma_{t}\mid\gamma\in\operatorname{Geo}(X),\gamma_{0}=x,\gamma_{1}\in A\right\}$ (13)
@@ 2.2. Quotients by isometric group actions
The action of $G$ on $X$ is proper if for any compact set $K\subseteq X$,
#$\#\{g\in G\mid gK\cap K\neq\emptyset\}<\infty,$ (17)
or, equivalently under our assumptions on $(X,\mathsf{d})$, that the map $(g,x)\mapsto(gx,x)$ is proper, see [Kap24, Thm. 11]. When the action of $G$ on $X$ is by isometries and proper, its orbits are closed, (16) is a distance, and $(X^{*},\mathsf{d}^{*})$ is a complete, separable, geodesic and proper metric space as well.
@@ 3.1. Sard properties and $\delta$-essentially non-branching
The *-minimizing Sard property was introduced in [BMR24, Def. 7.13], and it is a reinforcement of the minimizing Sard property. The “star” comes from the fact that, if the *-minimizing Sard property holds true, then for any $x\in M$ one can find a geodesically star-shaped subset where $\mathsf{d}^{2}(x,\cdot)$ is smooth, see [BMR24, Lemma 7.14] for a precise statement. If $M$ is a real-analytic manifold and the sub-Riemannian Hamiltonian is real-analytic, it is well-known that geodesics are either abnormal, or do not contain non-trivial abnormal segments; it follows that, in the real-analytic case, the *-minimizing and minimizing Sard properties are equivalent.
@@ 4. Discrete group actions
Let $(X,\mathsf{d},\mathsf{m})$ be a $\delta$-essentially non-branching metric measure space satisfying the $\mathrm{MCP}_{\mathrm{loc}}(K,N)$ for some $K\in\mathbb{R}$ and $N\in[1,\infty)$. Let $G$ be a discrete group of metric measure isometries of $(X,\mathsf{d},\mathsf{m})$ acting properly and freely. Then, the metric measure quotient $(X^{*},\mathsf{d}^{*},\mathsf{m}^{*})$ satisfies the $\mathrm{MCP}_{\mathrm{loc}}(K,N)$. Furthermore, if $(X,\mathsf{d},\mathsf{m})$ is $\delta$-essentially non-branching, then $(X^{*},\mathsf{d}^{*},\mathsf{m}^{*})$ is $\delta$-essentially non-branching.
@@
Let $D:=\mathrm{diam}(\operatorname{supp}(\mu_{0}^{*}),\operatorname{supp}(\mu_{1}^{*}))<\infty$. Consider the set of geodesics $\operatorname{supp}(\nu^{*})\subseteq\operatorname{Geo}(X^{*})$. The set of lifts $\theta(\operatorname{supp}(\nu^{*}))\subseteq\operatorname{Geo}(X)$ is contained in a compact ball of radius $D$ within the compact set $\operatorname{supp}(\mu_{0})$. Furthermore, the set of final points $e_{1}(\theta(\operatorname{supp}(\nu^{*})))$ is contained in the discrete set $\cup_{j=1}^{L^{*}}\mathsf{q}^{-1}(x_{j}^{*})$, which is then finite.
@@
By the $\delta$-essentially non-branching assumption, $\nu$ is concentrated on a set $\Gamma$ of non-branching geodesics, and we can assume that $\Gamma\subseteq\operatorname{supp}(\nu)\subseteq U$. The support of $\nu$ is made of geodesics that are lifts of geodesics in $\operatorname{supp}(\nu^{*})$, and two geodesics in $\operatorname{supp}(\nu)$ are branching on $X$ if and only if the corresponding projections are branching on $X^{*}$. It follows that $\mathsf{q}(\Gamma)\subseteq\operatorname{Geo}(X^{*})$ is a set of non-branching geodesics, and
@@ 5.3. Local measure contraction properties on the quotient
Let $(X,\mathsf{d},\mathsf{m})$ be a $\delta$-essentially non-branching metric measure space satisfying the $\mathrm{MCP}_{\mathrm{loc}}(K,N)$ for some $K\in\mathbb{R}$ and $N\in[1,\infty)$. Let $G$ be a connected, compact, abelian Lie group acting freely by metric measure isometries on $X$. Then, the metric measure quotient $(X^{*},\mathsf{d}^{*},\mathsf{m}^{*})$ satisfies the $\mathrm{MCP}_{\mathrm{loc}}(K,N)$.
@@ 6.9. The factorization argument
Let $\mathbb{M}=\mathbb{G}/\mathbb{H}$ be a Carnot homogeneous space, and let ~${\tilde{\mathbb{H}}}<\mathbb{G}$ be a closed, dilation-invariant subgroup with $\mathbb{H}\triangleleft{\tilde{\mathbb{H}}}$. Then the left multiplication of ${\tilde{\mathbb{H}}}$ on $\mathbb{G}$ induces a proper and free action of $R:={\tilde{\mathbb{H}}}/\mathbb{H}$ on $\mathbb{M}$ by metric measure isometries. The quotient by this action, equipped with the quotient metric, is smoothly isometric to the Carnot homogeneous space $\tilde{\mathbb{M}}=\mathbb{G}/{\tilde{\mathbb{H}}}$, in other words
@@
where, in the last equality, multiplication on the right by $g\in\mathbb{G}$ denotes the well-defined right action of $\mathbb{G}$ on $\mathbb{M}/\Gamma$, induced by the action on $\mathbb{M}$ (see Section 6.1), formally given by
@@
7. Measure contraction property on Carnot homogeneous spaces
@@
Suppose that a Carnot homogeneous space $\mathbb{M}=\mathbb{G}/\mathbb{H}$ admits a quotient to another Carnot homogeneous space $\tilde{\mathbb{M}}=\mathbb{G}/\tilde{\mathbb{H}}$. Assume that the minimizing Sard property holds for all factors $\mathbb{M},\mathbb{M}_{1},\dots,\mathbb{M}_{s-1}$ of the map $\pi:\mathbb{M}\to\tilde{\mathbb{M}}$ in Theorem 6.16 (this is the case, for example, if the minimizing Sard property holds for all Carnot homogeneous spaces of step $\leq s$, where $s$ is the step of $\mathbb{M}$). Then, if $\mathbb{M}$ satisfies the $\mathrm{MCP}(K,N)$ for some $K\in\mathbb{R}$ and $N\in[1,\infty)$, the space $\tilde{\mathbb{M}}$ also satisfies the $\mathrm{MCP}(K,N)$.
@@ 7.3. Carnot groups in low dimensions
Building on the previous results, we initiate the classification of Carnot groups that satisfy the $\mathrm{MCP}$. A Carnot group is indecomposable if its Lie algebra is not the direct sum of two non-trivial Lie algebras. A classification of indecomposable Carnot groups of dimension $\leq 7$ is provided in [LT22, Gon98], to which we refer for more details.
@@
Among the indecomposable Carnot groups of dimension $\leq 7$, listed in Table 1:
@@
Proof of Item i. The criterion for admitting a Martinet quotient of Theorem 6.13 is checked using the Lie algebra relations listed in [LT22] in terms of a specific basis. The Carnot groups marked in red in Table 1 are precisely all those that satisfy the criterion. A choice of subspaces $\mathfrak{h}_{3},\mathfrak{h}_{2}$ that verify the criterion is given in Table 2, and the computations are omitted.
@@
Table 2. Indecomposable Carnot groups of dimension $\leq 7$ admitting a quotient to Martinet, marked in red in Table 1. The generators for a choice of the subspaces $\mathfrak{h}_{2},\mathfrak{h}_{3}$ verifying Theorem 6.13 are given in terms of the bases of [LT22].
By Theorem 7.2, these Carnot groups – provided that the minimizing Sard property holds – do not satisfy any $\mathrm{MCP}(K,N)$. According to the discussion preceding Corollary 7.3, the Carnot groups for which the minimizing Sard property is known to be true are those of step $3$, all filiform ones (which, in Table 1, for step $\geq 4$, are $N_{5,2,1}$, $N_{6,2,1}$, and $123457A$), and Carnot groups of step $4$ and rank $2$ (which, in Table 1, are $N_{6,2,5}$, $N_{6,2,5a}$, $N_{6,2,7}$, and $2457L$). Otherwise, the failure of the $\mathrm{MCP}$ is conditional on the validity of the minimizing Sard property.
Proof of Item ii. In Carnot groups of step $1$ and $2$, it is well-known that all geodesics are normal, so the minimizing Sard property holds and there are no non-trivial Goh-Legendre geodesics. We then apply Corollary 7.5. All other Carnot groups marked in green in Table 1 are of step $3$. One can directly verify that all of them have a medium-fat distribution, and thus they do not have non-trivial Goh(-Legendre) abnormal geodesics (see [Rif14, Ex. 2.1]). Therefore, we can also apply Corollary 7.5 to them. The only exceptions are $N_{6,3,1}$, $247F_{1}$, $247P$ and $137A_{1}$, marked with a dagger in Table 1: their distribution is not medium-fat, and they have Goh geodesics, see Remark 7.7. Nevertheless, we can verify that they do not have non-trivial Goh–Legendre geodesics. For Carnot groups, the Goh–Legendre condition can be written in terms of the adjoint representation, and in step $3$, it admits the following simplified formulation (see Section A.1): there exists $\lambda_{0}\in\mathfrak{g}^{*}\setminus\{0\}$ such that
@@
The Goh condition alone is not sufficient in the proof. In fact, all Carnot groups marked with a dagger in Table 1 do have a non-trivial Goh geodesic. Such geodesics, however, do not satisfy the generalized Legendre condition. For $N_{6,3,1}$, $247F_{1}$, and $247P$, the geodesic is the one with control $u=(1,0,0)$. For $137A_{1}$, any horizontal curve with control $u=(0,0,u_{3},u_{4})$ satisfies the Goh condition.
@@
Item i follows from the tensorization property of the $\mathrm{MCP}$, see [Oht07b]. To prove Item ii, note that $\mathbb{G}$ admits a quotient to $\mathbb{G}_{i}$, and one can use Theorem 7.1. ∎
8. Failure of the measure contraction property for the Martinet structure
@@
$1+\frac{\omega\partial_{\omega}\mathcal{J}_{R}(\omega,m)}{\mathcal{J}_{R}(\omega,m)}\leq N$ (237)
will hold for all $m\in(0,1)$ and all $\omega\in(0,2\operatorname{K}(m)]$.
@@
By Theorem 8.4 if $\lambda_{0}\in T^{*}_{0}\mathbb{R}^{3}$ is such that $m\in(0,1)$ and $\omega\in(0,2\operatorname{K}(m)]$, then $\gamma:[0,1]\to\mathbb{R}^{3}$ defined by $\exp_{0}(t\lambda_{0})$ is a geodesic, $\lambda_{0}$ is not a critical point of $\exp_{0}$, and $\exp_{0}(\lambda_{0})\notin\mathrm{Cut}(0)$. Then, there is a neighborhood $\mathcal{U}$ of $\lambda_{0}$ such that $\exp_{0}:\mathcal{U}\to\exp_{0}(\mathcal{U})$ is a diffeomorphism, and such that for all $\lambda_{0}^{\prime}\in\mathcal{U}$, the geodesic $\gamma(t):=\exp_{0}(t\lambda_{0}^{\prime})$ is the unique geodesic joining 00 with $\exp_{0}(\lambda_{0}^{\prime})$ and $\exp_{0}(\lambda_{0}^{\prime})$ is not conjugate to 00 along $\gamma$. Assume that $\varepsilon>0$ is sufficiently small such that $B_{\varepsilon}(\exp_{0}(\lambda_{0}))\subseteq\exp_{0}(\mathcal{U})$ and let $A_{\varepsilon}$ be such that $\exp_{0}(A_{\varepsilon})=B_{\varepsilon}(\exp_{0}(\lambda_{0}))$. In particular by the $\mathrm{MCP}(0,N)$, we have that for all $t\in[0,1]$
@@
where $\mathscr{L}$ denotes the Lebesgue measure on $T_{0}^{*}\mathbb{R}^{3}\simeq\mathbb{R}^{3}$. From the expression of $\mathcal{J}^{t}$ we deduce
$|\mathcal{J}_{R}(\omega t,m)|\geq t^{N-1}|\mathcal{J}_{R}(\omega,m)|,$ (239)
for all $t\in[0,1]$, all $m\in(0,1)$ and all $\omega\in(0,2\operatorname{K}(m)]$. By the monotonicity, (239) is equivalent to
$\log|\mathcal{J}_{R}(\omega t,m)|-\log|\mathcal{J}_{R}(\omega,m)|\geq(N-1)\log t.$ (240)
This last inequality can be rewritten using the fundamental theorem of calculus as
$\int^{\omega}_{\omega t}\frac{\mathrm{d}}{\mathrm{d}z}\log|\mathcal{J}_{R}(z,m)|\mathrm{d}z\leq(N-1)\int^{\omega}_{\omega t}\frac{\mathrm{d}}{\mathrm{d}z}\log z\,\mathrm{d}z.$ (241)
Since this inequality has to hold for all $\omega\in(0,2\operatorname{K}(m)]$ and all $t\in[0,1]$, it is equivalent to the same inequality on the integrands:
$\frac{\partial_{\omega}\mathcal{J}_{R}(\omega,m)}{\mathcal{J}_{R}(\omega,m)}\leq(N-1)\frac{1}{\omega},$ (242)
for all $m\in(0,1)$ and $\omega\in(0,2\operatorname{K}(m)]$. ∎
The failure of the measure contraction property can finally be obtained by showing that the differential inequality (237) does not hold in the Martinet structure.

### v2

@@ 2.1. Measure contraction property
The $\mathrm{MCP}$ implies a corresponding Bishop-Gromov volume growth inequality for $(X,\mathsf{d},\mathsf{m})$, and all its consequences. In particular, we record that a metric measure space satisfying the $\mathrm{MCP}(K,N)$ with $K>0$ (and $N>1$) must be compact, and that setting
$Z_{t}(x,A):=\left\{\gamma_{t}\mid\gamma\in\operatorname{Geo}(X),\gamma_{0}=x,\gamma_{1}\in A\right\}$ (15)
@@ 2.2. Quotients by isometric group actions
The action of $G$ on $X$ is proper if the map $(g,x)\mapsto(gx,x)$ is proper. When the action of $G$ on $X$ is by isometries and proper, its orbits are closed, (18) is a distance, and $(X^{*},\mathsf{d}^{*})$ is a complete, separable, geodesic and proper metric space as well.
@@ 3.1. Sard properties and $\delta$-essentially non-branching
The *-minimizing Sard property was introduced in [BMR24, Def. 7.13], and it is a reinforcement of the minimizing Sard property. The “star” comes from the fact that, if the *-minimizing Sard property holds true, then for any $x\in M$ one can find a geodesically star-shaped and full-measure subset where $\mathsf{d}^{2}(x,\cdot)$ is smooth, see [BMR24, Lemma 7.14]. If $M$ is a real-analytic manifold and the sub-Riemannian Hamiltonian is real-analytic, it is well-known that geodesics are either abnormal, or do not contain non-trivial abnormal segments; it follows that, in the real-analytic case, the *-minimizing and minimizing Sard properties are equivalent.
@@ 4. Discrete group actions
Let $(X,\mathsf{d},\mathsf{m})$ be a metric measure space satisfying the $\mathrm{MCP}_{\mathrm{loc}}(K,N)$ for some $K\in\mathbb{R}$ and $N\in[1,\infty)$. Let $G$ be a discrete group of metric measure isometries, acting properly and freely. Then, the metric measure quotient $(X^{*},\mathsf{d}^{*},\mathsf{m}^{*})$ satisfies the $\mathrm{MCP}_{\mathrm{loc}}(K,N)$. If $(X,\mathsf{d},\mathsf{m})$ is $\delta$-essentially non-branching, then $(X^{*},\mathsf{d}^{*},\mathsf{m}^{*})$ is $\delta$-essentially non-branching.
@@
Let $D:=\mathrm{diam}(\operatorname{supp}(\mu_{0}^{*}),\operatorname{supp}(\mu_{1}^{*}))<\infty$. Consider the set of geodesics $\operatorname{supp}(\nu^{*})\subseteq\operatorname{Geo}(X^{*})$. The set of lifts $\theta(\operatorname{supp}(\nu^{*}))\subseteq\operatorname{Geo}(X)$ is contained in a compact ball of radius $D$ within the compact set $\operatorname{supp}(\mu_{0})$. Furthermore, the set of final points $e_{1}(\theta(\operatorname{supp}(\nu^{*})))$ is contained in the discrete set $\cup_{j=1}^{L^{*}}\mathsf{q}^{-1}(x_{j}^{*})$, so that $e_{1}(\theta(\operatorname{supp}(\nu^{*})))$ is finite.
@@
By the $\delta$-essentially non-branching assumption, $\nu$ is concentrated on a set $\Gamma$ of non-branching geodesics, and we can assume that $\Gamma\subseteq\operatorname{supp}(\nu)$. The support of $\nu$ is made of geodesics that are lifts of geodesics in $\operatorname{supp}(\nu^{*})$, and two geodesics in $\operatorname{supp}(\nu)$ are branching on $X$ if and only if the corresponding projections are branching on $X^{*}$. It follows that $\mathsf{q}(\Gamma)\subseteq\operatorname{Geo}(X^{*})$ is a set of non-branching geodesics, and
@@ 5.3. Local measure contraction properties on the quotient
Let $(X,\mathsf{d},\mathsf{m})$ be a $\delta$-essentially non-branching metric measure space satisfying the $\mathrm{MCP}_{\mathrm{loc}}(K,N)$ for some $K\in\mathbb{R}$ and $N\in[1,\infty)$. Let $G$ be a connected, compact, abelian Lie group of metric measure isometries, acting freely. Then, the metric measure quotient $(X^{*},\mathsf{d}^{*},\mathsf{m}^{*})$ satisfies the $\mathrm{MCP}_{\mathrm{loc}}(K,N)$.
@@ 6.9. The factorization argument
Let $\mathbb{M}=\mathbb{G}/\mathbb{H}$ be a Carnot homogeneous space, and let ~${\tilde{\mathbb{H}}}<\mathbb{G}$ be a closed, dilation-invariant subgroup with $\mathbb{H}\triangleleft{\tilde{\mathbb{H}}}$. Then the left multiplication of ${\tilde{\mathbb{H}}}$ on $\mathbb{G}$ induces a proper and free action of $R:={\tilde{\mathbb{H}}}/\mathbb{H}$ on $\mathbb{M}$ by metric measure isometries. The quotient of $\mathbb{M}$ by this action, equipped with the quotient metric, is smoothly isometric to the Carnot homogeneous space $\tilde{\mathbb{M}}=\mathbb{G}/{\tilde{\mathbb{H}}}$, in other words
@@
where, in the last equality, the multiplication on the right by $g\in\mathbb{G}$ denotes the well-defined right action of $\mathbb{G}$ on $\mathbb{M}/\Gamma$, induced by the action on $\mathbb{M}$ (see Section 6.1), formally given by
@@
7. $\mathrm{MCP}$ on Carnot homogeneous spaces
@@
Suppose that a Carnot homogeneous space $\mathbb{M}=\mathbb{G}/\mathbb{H}$ admits a quotient to a Carnot homogeneous space $\tilde{\mathbb{M}}=\mathbb{G}/\tilde{\mathbb{H}}$. Assume that the minimizing Sard property holds for all factors $\mathbb{M},\mathbb{M}_{1},\dots,\mathbb{M}_{s-1}$ of the map $\pi:\mathbb{M}\to\tilde{\mathbb{M}}$ in Theorem 6.16 (this is the case, for example, if the minimizing Sard property holds for all Carnot homogeneous spaces of step $\leq s$, where $s$ is the step of $\mathbb{M}$). Then, if $\mathbb{M}$ satisfies the $\mathrm{MCP}(K,N)$ for some $K\in\mathbb{R}$ and $N\in[1,\infty)$, the space $\tilde{\mathbb{M}}$ also satisfies the $\mathrm{MCP}(K,N)$.
@@ 7.3. Carnot groups in low dimensions
Building on the previous results, we initiate the classification of Carnot groups that satisfy the $\mathrm{MCP}$. A stratified group is indecomposable if its Lie algebra is not the direct sum of two non-trivial Lie algebras. A classification of indecomposable stratified groups of dimension $\leq 7$ is provided in [LT22, Gon98], to which we refer for details.
@@
Among the indecomposable stratified groups of dimension $\leq 7$, listed in Table 1, equipped with any choice of Carnot-Carathéodory metric and invariant measure turning it into a Carnot group:
@@
Proof of Item i. The criterion for admitting a Martinet quotient of Theorem 6.13 is checked using the Lie algebra relations listed in [LT22] in terms of a specific basis. The groups marked in red in Table 1 are precisely all those that satisfy the criterion. A choice of subspaces $\mathfrak{h}_{3},\mathfrak{h}_{2}$ that verify the criterion is given in Table 2, and the computations are omitted.
@@
Table 2. Indecomposable stratified groups of dimension $\leq 7$ admitting a quotient to Martinet, marked in red in Table 1. The generators for a choice of the subspaces $\mathfrak{h}_{2},\mathfrak{h}_{3}$ verifying Theorem 6.13 are given in terms of the bases of [LT22].
Choose a scalar product on the first layer giving these stratified groups the structure of a Carnot group. By Theorem 7.2, these Carnot groups – provided that the minimizing Sard property holds – do not satisfy any $\mathrm{MCP}(K,N)$. According to the discussion preceding Corollary 7.3, the Carnot groups for which the minimizing Sard property is known to be true are those of step $3$, all filiform ones (which, in Table 1, for step $\geq 4$, are $N_{5,2,1}$, $N_{6,2,1}$, and $123457A$), and Carnot groups of step $4$ and rank $2$ (which, in Table 1, are $N_{6,2,5}$, $N_{6,2,5a}$, $N_{6,2,7}$, and $2457L$).
Proof of Item ii. In Carnot groups of step $1$ and $2$, it is well-known that all geodesics are normal, so the minimizing Sard property holds and there are no non-trivial Goh-Legendre geodesics. We then apply Corollary 7.5. All other groups marked in green in Table 1 are of step $3$. One can directly verify that all of them have a medium-fat distribution, and thus they do not have non-trivial Goh(-Legendre) abnormal horizontal curves (see [Rif14, Ex. 2.1]). Therefore, we can also apply Corollary 7.5 to them. The only exceptions are $N_{6,3,1}$, $247F_{1}$, $247P$ and $137A_{1}$, marked with a dagger in Table 1: their distribution is not medium-fat, and they have Goh geodesics, see Remark 7.7. Nevertheless, we can verify that they do not have non-trivial Goh–Legendre geodesics. For stratified groups, the Goh–Legendre condition can be written in terms of the adjoint representation, and in step $3$, it admits the following simplified formulation (see Section A.1): there exists $\lambda_{0}\in\mathfrak{g}^{*}\setminus\{0\}$ such that
@@
The Goh condition alone is not sufficient in the proof. In fact, all groups marked with a dagger in Table 1 do have a non-trivial Goh horizontal curves. Such curves, however, do not satisfy the generalized Legendre condition. For $N_{6,3,1}$, $247F_{1}$, and $247P$, such curves have control $u=(u_{1},0,0)$. For $137A_{1}$, any horizontal curve with control $u=(0,0,u_{3},u_{4})$ satisfies the Goh condition. Taking these controls to be constant corresponds to segments in the first stratum, that are geodesics for any choice of sub-Riemannian metric.
@@
Item i follows from the well-known tensorization property of the $\mathrm{MCP}$, see [Oht07b]. To prove Item ii, note that $\mathbb{G}$ admits a quotient to $\mathbb{G}_{i}$, and one can use Theorem 7.1. ∎
8. Failure of the $\mathrm{MCP}$ for generic structures
Purpose of this section is to prove Theorem 1.19, which we recall here.
Theorem 8.1 (Generic failure of the $\mathrm{MCP}$ for rank $3$ and high dimension).
Assume that the minimizing Sard property holds for Carnot groups. In the same setting of LABEL:{gen:genideal}, if
$\dim M\geq(k-1)\left(\frac{k^{2}}{3}+\frac{5k}{6}+1\right),$ (216)
then there exists an open dense subset $W_{k}^{\prime}\subseteq W_{k}$ of $\mathcal{G}_{k}$ such that for every element of $W_{k}^{\prime}$, the corresponding sub-Riemannian structure does not admit non-trivial abnormal geodesics and does not satisfy the $\mathrm{MCP}(K,N)$ for any $K\in\mathbb{R}$, $N\in[1,\infty)$ and any choice of smooth measure.
Proof.
Remember the set $W_{k}$ from Theorem 1.18. By [AG01, Lemma 4], if (216) holds, then the growth vector at $x$, i.e. $(\dim\mathscr{D}_{x}^{1},\dim\mathscr{D}_{x}^{2},\dots,\dim\mathscr{D}_{x}^{s(x)})$, is generically maximal, and the tangent cone of $(M,\mathsf{d})$ at $x$ admits a quotient to the Martinet structure. This means that, up to taking a smaller open dense set $W_{k}^{\prime}\subseteq W_{k}$, for any pair $(\mathscr{D},g)\in W_{k}^{\prime}$ we can apply Theorem 7.2. ∎
We also prove Theorem 1.21, whose statement we recall here.
Theorem 8.2 (Failure of the $\mathrm{MCP}$ for rank $2$).
Let $(M,\mathsf{d},\mathsf{m})$ be a sub-Riemannian metric measure space with constant rank $k=2$. Assume that there exists $x\in M$ such that
$\dim\mathscr{D}^{3}_{x}-\dim\mathscr{D}^{2}_{x}\geq 1.$ (217)
Assume that the minimizing Sard property holds for all Carnot groups of step $\leq s(x)$. Then $(M,\mathsf{d},\mathsf{m})$ does not satisfy the $\mathrm{MCP}(K,N)$ for any $K\in\mathbb{R}$ and $N\in[1,\infty)$.
Proof.
Under these assumptions, the growth vector at $x$ is either $(2,2,3,\dots)$, $(2,2,4,\dots)$, $(2,3,4,\dots)$, or $(2,3,5,\dots)$. In each case the tangent cone of $(M,\mathsf{d})$ at $x$ admits a quotient to the Martinet structure (see also the proof of [AG01, Prop. 7]). We conclude using Theorem 7.2. ∎
9. Failure of the $\mathrm{MCP}$ for the Martinet structure
@@
$|\mathcal{J}_{R}(\omega t,m)|\geq t^{N-1}|\mathcal{J}_{R}(\omega,m)|$ (240)
will hold for all $t\in[0,1]$, all $m\in(0,1)$ and all $\omega\in(0,2\operatorname{K}(m))$.
@@
By Theorem 9.4 if $\lambda_{0}\in T^{*}_{0}\mathbb{R}^{3}$ is such that $m\in(0,1)$ and $\omega\in(0,2\operatorname{K}(m))$, then $\gamma:[0,1]\to\mathbb{R}^{3}$ defined by $\exp_{0}(t\lambda_{0})$ is a geodesic, $\lambda_{0}$ is not a critical point of $\exp_{0}$, and $\exp_{0}(\lambda_{0})\notin\mathrm{Cut}(0)$. Then, there is a neighborhood $\mathcal{U}$ of $\lambda_{0}$ such that $\exp_{0}:\mathcal{U}\to\exp_{0}(\mathcal{U})$ is a diffeomorphism, and such that for all $\lambda_{0}^{\prime}\in\mathcal{U}$, the geodesic $\gamma(t):=\exp_{0}(t\lambda_{0}^{\prime})$ is the unique geodesic joining 00 with $\exp_{0}(\lambda_{0}^{\prime})$ and $\exp_{0}(\lambda_{0}^{\prime})$ is not conjugate to 00 along $\gamma$. Assume that $\varepsilon>0$ is sufficiently small such that $B_{\varepsilon}(\exp_{0}(\lambda_{0}))\subseteq\exp_{0}(\mathcal{U})$ and let $A_{\varepsilon}$ be such that $\exp_{0}(A_{\varepsilon})=B_{\varepsilon}(\exp_{0}(\lambda_{0}))$. In particular by the $\mathrm{MCP}(0,N)$ and Eq. 16 in particular, we have that for all $t\in[0,1]$
@@
where $\mathscr{L}$ is the Lebesgue measure on $T_{0}^{*}\mathbb{R}^{3}\simeq\mathbb{R}^{3}$. We deduce (240) with the expression (233). ∎
The failure of the measure contraction property can finally be obtained by showing that (240) does not hold in the Martinet structure.